# Amplifier: Bringing Attention to Neglected Low-Energy Components in Time Series Forecasting

### v1

@@ 1 Introduction
Low-energy components are indispensable. In contexts like signal compression or model simplification, low-energy components might be dismissed as unimportant or treated as noise. However, in scenarios where subtle changes or background patterns are significant, low-energy components can hold critical details. For example, in weather forecasting, small meteorological shifts can accumulate over time, leading to dramatic changes in weather patterns. In financial markets, minor fluctuations can trigger larger trends or volatility, with high-frequency trading systems often capitalizing on these small changes to execute profitable trades. We also have empirically validated the indispensability of low-energy components in time series forecasting. As illustrated in Figure 1(a), directly filtering out low-energy components (as shown in the right figure of (a)) leads to increased MSE values across different types of models (PatchTST, RLinear, DLinear, FreTS), highlighting the crucial role of low-energy components in enhancing prediction accuracy.
Low-energy components are often overlooked. We generate two signals with different energy distributions (corresponding to the top and bottom of Figure 1(b)), and conduct iTransformer (Liu et al. 2024), PatchTST (Nie et al. 2023), and RLinear (Li et al. 2023) on them respectively. As detailed in Figure 1(b), the phenomenon of ignorance (rf. small green circles) always occurs with low-energy components, regardless of the frequency band in which they are located.
@@ 2 Related Work
2.1 Deep Time Series Forecasting
@@
2.2 Advancements in Frequency Domain Techniques
@@ 3 Preliminaries
Time Series Forecasting.
@@
The proof is shown in Appendix A. It indicates that during the training process, the parameters associated with low-energy components update more slowly and less efficiently than those associated with high-energy components. As a result, by the end of training, the parameters associated with low-energy components are further from their true values compared to those associated with high-energy components, indicating that these parameters have minimal participation during training, making them appear “ignored”.
@@ 4 Methodology
4.1 Energy Amplification Technique
@@ 5.3 Model Analysis
Dataset ETTh1 ETTm2 Weather
Metrics MSE MAE MSE MAE MSE MAE
w/o EAT 0.393 0.407 0.182 0.266 0.184 0.230
Amplifier 0.371 0.392 0.176 0.258 0.156 0.205
Boost 5.643% 3.613% 3.484% 2.973% 15.100% 10.923%
(a) iTransformer
(b) Autoformer
(c) SparseTSF
(d) DLinear
@@
5.4 Efficiency Analysis
The theoretical complexity of the Amplifier is $\mathcal{O}(L\log L)$. We conduct a comprehensive comparison of the forecasting performance, the scale of parameters, and the training speed of the following representative models, including SparseTSF, RLinear, DLinear, FITS, FreTS, and iTransformer. We choose the ETTh1 and Weather datasets under the scenario of $L=96$ and $\tau=96$ for comparison.
@@ 5.5 Visualizations
(a) Amplifier
(b) FreTS
(c) iTransformer
(d) DLinear
@@
Visualization of Forecasting Results
To visually compare the performance of Amplifier with state-of-the-art models, including iTransformer, FreTS, and DLinear, we present prediction showcases on the ETTm2 and Electricity datasets, and the results are shown in Figures 6 and 7. Compared to these different types of state-of-the-art models, Amplifier provides the most accurate predictions of future series variations, demonstrating superior performance.

### v3

@@ 1 Introduction
Low-energy components are indispensable. In contexts like signal compression or model simplification, low-energy components might be dismissed as unimportant or treated as noise. However, in scenarios where subtle changes or background patterns are significant, low-energy components can hold critical details. For example, in weather forecasting, small meteorological shifts can accumulate over time, leading to dramatic changes in weather patterns. In financial markets, minor fluctuations can trigger larger trends or volatility, with high-frequency trading systems often capitalizing on these small changes to execute profitable trades. We also have empirically validated the indispensability of low-energy components in time series forecasting. As illustrated in Figure 1(a), directly filtering out low-energy components leads to increased MSE values across different types of models (PatchTST, RLinear, DLinear, FreTS), highlighting the crucial role of low-energy components in enhancing prediction accuracy.
Low-energy components are often overlooked. We generate two signals with different energy distributions (corresponding to the left and right of Figure 1(b)), and conduct iTransformer (Liu et al. 2024), PatchTST (Nie et al. 2023), and DLinear (Zeng et al. 2023) on them respectively. As detailed in Figure 1(b), the phenomenon of ignorance (rf. small green circles) always occurs with low-energy components, regardless of the frequency band in which they are located.
@@ 2 Related Work
Deep Time Series Forecasting
@@
Advancements in Frequency Domain Techniques
@@ 3 Preliminaries
Time Series Forecasting
@@
The proof is shown in Appendix A. It indicates that during the training process, the parameters associated with low-energy components update more slowly and less efficiently than those associated with high-energy components. As a result, by the end of training, the parameters associated with low-energy components are further from their true values compared to those associated with high-energy components, indicating that these parameters have minimal participation during training, making them appear ”ignored”.
@@ 4 Methodology
Energy Amplification Technique
@@ 5 Experiments
(a) iTransformer
(b) Autoformer
(c) SparseTSF
(d) DLinear
Dataset ETTh1 ETTm2 Weather
Metrics MSE MAE MSE MAE MSE MAE
w/o EAT 0.393 0.407 0.182 0.266 0.184 0.230
Amplifier 0.371 0.392 0.176 0.258 0.156 0.205
Boost 5.643% 3.613% 3.484% 2.973% 15.100% 10.923%
@@
Efficiency Analysis
The theoretical complexity of the Amplifier is $\mathcal{O}(L\log L)$. We conduct a comprehensive comparison of the forecasting performance, the scale of parameters, and the training speed of the following representative models, including SparseTSF, RLinear, DLinear, FITS, FreTS, and iTransformer. We choose the Weather dataset under the scenario of $L=96$ and $\tau=96$ for comparison.
@@
Visualization of Forecasting Results
To visually compare the performance of Amplifier with state-of-the-art models, including iTransformer, FreTS, and DLinear, we present prediction showcases on the ETTm2 and Electricity datasets, and the results are shown in Figures 6 and 7. The red line represents the input sequence, the blue line represents the ground truth, and the yellow line represents the predicted value. Compared to these different types of state-of-the-art models, Amplifier provides the most accurate predictions of future series variations, demonstrating superior performance.